# NeuralOS: Towards Simulating Operating Systems via Neural Generative Models

### v1

@@ Abstract
We introduce NeuralOS, a neural framework that simulates graphical user interfaces (GUIs) of operating systems by directly predicting screen frames in response to user inputs such as mouse movements, clicks, and keyboard events. NeuralOS combines a recurrent neural network (RNN), which tracks computer state, with a diffusion-based neural renderer that generates screen images. The model is trained on a large-scale dataset of Ubuntu XFCE recordings, which include both randomly generated interactions and realistic interactions produced by AI agents. Experiments show that NeuralOS successfully renders realistic GUI sequences, accurately captures mouse interactions, and reliably predicts state transitions like application launches. Although modeling fine-grained keyboard interactions precisely remains challenging, NeuralOS offers a step toward creating fully adaptive, generative neural interfaces for future human-computer interaction systems.
@@ 1 Introduction
Recent breakthroughs in generative models have transformed human-computer interaction, making it increasingly adaptive, personalized, and intuitive. Historically, computing interfaces were rigid and predefined, such as command-line terminals and static graphical menus (Engelbart, 1968). The emergence of large language models (LLMs) and multimodal AI systems expanded this paradigm by enabling interactions through natural language (Radford et al., 2019; Brown et al., 2020), images (Ho et al., 2020; Lipman et al., 2022; Radford et al., 2021; Song et al., 2020b), and videos (OpenAI, 2024). Recently, generative models have even begun simulating dynamic visual environments (Ha and Schmidhuber, 2018a; He et al., 2025), notably interactive video games (Alonso et al., 2024; Feng et al., 2024; Oh et al., 2015; Valevski et al., 2024). These advancements suggest a future where computing interfaces could become fully generative, dynamically adapting in real-time based on user inputs, contexts, and intentions (Deka et al., 2017).
In this paper, we introduce NeuralOS, a first step toward realizing this vision. NeuralOS simulates an operating system’s graphical interface entirely using deep neural networks. By modeling the OS interface as a generative process, it directly predicts graphical frames from user input events, such as mouse movements, clicks, and keyboard interactions, without manually programmed kernels or applications. Figure 1 illustrates an example sequence generated by NeuralOS, demonstrating realistic cursor movements and window interactions predicted solely from user inputs.
Figure 1: Example image sequence predicted by NeuralOS, illustrating the model’s ability to simulate realistic GUI interactions. The sequence shows key frames as a user: (a-b) moves the cursor to the “Home” icon, (c-d) double-clicks to open the “Home” folder, (e) moves the cursor toward the window’s close button, and (f) clicks to close the window. Cursor positions are highlighted with red circles. Frames are generated autoregressively, conditioned on previous frames and user inputs.
NeuralOS integrates two complementary neural architectures, analogous to the traditional separation between OS kernels and desktop rendering programs: a recurrent neural network (RNN) (Hochreiter and Schmidhuber, 1997) maintains internal computer states (such as open applications, hidden windows, and recent actions), while a diffusion-based convolutional neural renderer generates screen images. We train NeuralOS end-to-end on interaction sequences recorded from Ubuntu XFCE environments, combining randomly generated and realistic AI-generated human-like interactions.
@@
Experiments show that NeuralOS can generate realistic screen sequences, accurately predict mouse interactions, and reliably simulate transitions such as application openings. While computational constraints limit our model’s ability to capture fine-grained keyboard inputs precisely, NeuralOS offers a step toward neural operating systems that adapt interfaces in real time, potentially enabling users to personalize interactions through natural language or gestures instead of fixed menus. More broadly, this work suggests the exciting possibility of blurring boundaries between software applications. For example, in the future, passive media such as movies could be transformed into interactive experiences (Yu et al., 2025). These results point toward user interfaces that may eventually become fully AI-driven, highly flexible, and tailored to individual needs. Our code, pretrained model, and an interactive demo are available at https://neural-os.com.
@@ 2 Generative Modeling of Operating System Interfaces
Formally, each frame $x_{t}$ is represented as an image tensor $x_{t}\in\mathbb{R}^{H\times W\times C}$, with $H$ and $W$ denoting image height and width, respectively, and $C$ the number of color or feature channels. The input event $a_{t}$ at timestep $t$ includes cursor coordinates $(x,y)$, binary indicators for mouse clicks (left or right), and a binary vector indicating which keyboard keys are pressed or released. [^1] We assume user inputs are discretized and associated with each graphical frame, aggregating any events occurring between frames at the subsequent timestep.
The joint probability distribution of an OS graphical sequence given user inputs can be expressed as:
@@
where $\theta$ represents the parameters of our neural generative model.
Unlike standard video generation, OS simulation must respond instantly to unpredictable user inputs, often causing abrupt changes in the interface, such as when a new application is launched. This contrasts with the smooth, predictable transitions typical in video generation. As a result, the model must maintain accurate and responsive state tracking. Next, we describe the NeuralOS architecture and training strategies designed for these requirements.
Time Step 1Time Step 2Time Step $t$ Frame $x_{1}$ Renderer Renderer Context RNNRefer to captionRefer to caption Frame $x_{2}$ Renderer Renderer Context RNNRefer to captionRefer to $\cdots$ Frame $x_{t}$ Renderer Renderer Context RNNRefer to captionRefer to caption
@@ 3 NeuralOS Architecture
NeuralOS employs a hierarchical two-level RNN architecture to track the system state (see Figure 2(b)). Unlike transformers (Vaswani et al., 2017), whose inference complexity increases with context length, the RNN maintains constant complexity per timestep, which is crucial for continuous, long-horizon OS simulation.
At each timestep $t$, user inputs $a_{t}$ are encoded into embeddings. Specifically, cursor coordinates are discretized screen positions ($a_{t}^{x}$, $a_{t}^{y}$), mouse clicks are binary indicators, and keyboard keys are binary press/release states. Each component is embedded separately and then concatenated: [^2] We initialize keyboard embeddings such that the “release” state corresponds to the zero vector.
@@
To ensure that the lower-level LSTM maintains awareness of higher-level contextual information, the upper-level LSTM’s output $U_{t}$ is fed back as an input to the lower-level LSTM in the next timestep.
@@
Precise cursor localization is critical for realistic OS interactions. NeuralOS explicitly encodes cursor positions using a Gaussian spatial map $E_{\text{pos}}=M_{t}\in\mathbb{R}^{H\times W}$. Instead of using a one-hot cursor position (which can lose precision due to latent resolution constraints), we construct a Gaussian map centered at the cursor’s scaled coordinates:
@@
As demonstrated in Section 7, this spatial encoding is vital for accurate cursor rendering. This map $M_{t}$, combined with LSTM outputs $L_{t}$, $U_{t}$, forms the renderer context $R_{t}\in\mathbb{R}^{H\times W\times C^{\prime}}$:
@@
We concatenate the noisy image with $R_{t}$ as input to the UNet, and then predict the clean image. [^3] While most diffusion models are trained to predict noise (Ho et al., 2020), we find that predicting the clean image yields better performance in our setting. This may be because our RNN is pretrained to output the clean image, which is then fed into the diffusion UNet. Using a matching target for the UNet may better preserve the signal from the RNN output.
@@
Figure 3: Multi-stage training pipeline for NeuralOS. (1) RNN Pretraining: The RNN is pretrained to predict latent frames using a mean squared error (MSE) loss. (2) Joint Training: The pretrained RNN and the diffusion-based renderer are jointly optimized using a standard diffusion loss. (3) Scheduled Sampling: To mitigate error accumulation caused by exposure bias, the most recent input frame is occasionally replaced by a previously generated frame during training. (4) Context Length Extension: Input context is extended to enable the model to capture long-term dependencies.
@@ 4 Multi-Stage Training Approach
Training NeuralOS is practically challenging due to ineffective use of RNN outputs, error accumulation during inference, and difficulties in capturing long-term dependencies due to computational constraints. To address these challenges, we take a multi-stage training approach (Figure 3).
@@
Unlike text-to-image diffusion models such as Stable Diffusion (Rombach et al., 2022), which use pretrained textual encoders, NeuralOS uses a customized RNN without pretrained checkpoints. Our initial experiments show that direct joint training often leads to the renderer ignoring RNN outputs, as indicated by negligible gradient flow into the RNN. The diffusion-based renderer receives two streams of inputs: noisy latent frames and the RNN output; and without proper initialization of the RNN, it relies solely on the noisy image inputs, resulting in an under-trained RNN.
To address this, we first pretrain the RNN. Specifically, we structure the RNN output $R_{t}\in\mathbb{R}^{H\times W\times C^{\prime}}$ to match the spatial dimensions of the latent frame $x_{t}\in\mathbb{R}^{H\times W\times C}$, but with more channels ($C^{\prime}>C$). The RNN is pretrained to predict the latent frames $x_{t}$ using a mean squared error (MSE) loss:
@@
We jointly optimize the pretrained RNN and the diffusion-based renderer with a standard diffusion loss (Ho et al., 2020). The meaningful latent representations learned in RNN pretraining enable the renderer to use the RNN outputs, thus preventing the RNN outputs from being disregarded.
@@
During inference, the errors accumulate over time progressively degrade the quality of the generated frames. This issue arises from exposure bias (Ranzato et al., 2015): a model trained exclusively on ground-truth frames becomes overly reliant on perfect inputs and struggles when forced to operate on its own imperfect predictions during inference.
@@
Although NeuralOS can model sequences of arbitrary length, hardware memory limits require shorter sequences during training, which limits exposure to long-term dependencies. To address this, we introduce a final stage that extends training to longer contexts, following initial training on short context windows for efficiency.
To help the model distinguish true sequence beginnings from truncated ones, we use two distinct learnable initial states for the RNN: one for genuine starts and one for mid-sequence truncations. This distinction allows the model to manage uncertainties, depending on whether it observes the full input sequence or begins mid-interaction.
@@ 5 Data Collection
To collect realistic user interactions, we use Anthropic’s Claude-3.5-Sonnet computer-use agent (Anthropic, 2024), which processes screenshots and invokes provided interaction functions. To maximize interaction diversity, we structure the agent’s exploration process around a state-space search tree representing various OS states (Figure 4).
@@
In generating these random interactions, we simulate mouse movements, clicks, and keyboard inputs (key presses and releases) stochastically. To improve realism, we introduce several constraints and heuristics iteratively developed through experimentation. Cursor movements are modeled using Bezier curves to emulate natural mouse trajectories. Double-click events, rare under purely random sampling, are explicitly generated. Additionally, we enforce constraints such as limiting simultaneous key presses and ensuring keys are released only if previously pressed. Detailed implementation specifics are provided in our open-source codebase.
Large-Scale Data Collection Infrastructure
For efficient and scalable data collection, we use Docker containers that support parallel data collection. To simplify the dataset and ensure feasible model training, we customize the desktop environment with a minimal set of applications and a relatively low screen resolution ($512\times 384$).
Figure 5: NeuralOS Evaluation Results. (a) Heatmap illustrating predicted vs. ground truth state transitions. Each cell represents the percentage of predictions assigned to a particular predicted cluster (x-axis), given a ground-truth cluster (y-axis). Only the top 16 clusters are displayed here; refer to Figure 9 for the complete heatmap. (b) Comparison of cursor position errors for NeuralOS (with cursor position map), NeuralOS without the cursor position map, and a random baseline.
@@ 6 Experimental Setup
We collected data using 64 parallel Docker containers, each configured with Ubuntu 20.04 and XFCE desktops at a resolution of $512\times 384$. Data consisted of 2K agent-based and 120K random exploration demonstrations, each 30 seconds long at 15 frames per second (fps), resulting in roughly 12TB of latent data after compression via an autoencoder. The autoencoder reduced the images by a factor of 8 to a latent resolution of $64\times 48$ with 16 channels per frame. Detailed hyperparameters of the autoencoder are in Appendix D. A subset of 40K videos was reserved for evaluation.
Model Architecture
The hierarchical RNN has two LSTM modules (each with hidden size 4,096) and a multi-headed attention module (8 heads, 1,024 total dimension). The RNN output is projected to 32 channels and concatenated with the noisy latent frame (16 channels), resulting in a 48-channel input to the UNet renderer. The UNet uses four resolution levels with channel multipliers of [1, 2, 3, 5], two residual blocks per level, and attention layers at resolutions 8, 4, and 2. It has a base model dimension of 192 and outputs 16 channels. The final model contains 2.2B parameters for the RNN and 263M parameters for the renderer.
@@
NeuralOS was trained using our proposed multi-stage training approach. See Appendix E for full details. The total data processing and training took approximately 4 months, requiring 17,000 GPU hours on a server with 8 NVIDIA H200 GPUs (141GB memory per GPU) and an additional 6,000 GPU hours on a server with 8 NVIDIA H100 GPUs (80GB memory per GPU). At inference time, we used DDIM sampling (Song et al., 2020a) with 32 steps, achieving an inference speed of 1.8 fps on a single NVIDIA H100 GPU.

### v2

@@ Abstract
We introduce NeuralOS, a neural framework that simulates graphical user interfaces (GUIs) of operating systems by directly predicting screen frames in response to user inputs such as mouse movements, clicks, and keyboard events. NeuralOS combines a recurrent neural network (RNN), which tracks computer state, with a diffusion-based neural renderer that generates screen images. The model is trained on a dataset of Ubuntu XFCE recordings, which include both randomly generated interactions and realistic interactions produced by AI agents. Experiments show that NeuralOS successfully renders realistic GUI sequences, accurately captures mouse interactions, and reliably predicts state transitions like application launches. Beyond reproducing existing systems, NeuralOS shows that synthesized training data can teach the model to simulate applications that were never installed, as illustrated by a Doom application, and suggests a path toward learning user interfaces purely from synthetic demonstrations.
@@ 1 Introduction
Recent breakthroughs in generative models have transformed human-computer interaction, making it increasingly adaptive, personalized, and intuitive. Historically, computing interfaces were rigid and predefined, such as command-line terminals and static graphical menus (Engelbart, 1968). The emergence of large language models (LLMs) and multimodal AI systems expanded this paradigm by enabling interactions through natural language (Radford et al., 2019; Brown et al., 2020), images (Ho et al., 2020; Lipman et al., 2022; Radford et al., 2021; Song et al., 2020b), and videos (Ho et al., 2022b; Singer et al., 2022b; OpenAI, 2024). Recently, generative models have even begun simulating dynamic visual environments (Ha & Schmidhuber, 2018a; He et al., 2025a), notably interactive video games (Feng et al., 2024; Oh et al., 2015; Valevski et al., 2024). These advancements suggest a future where computing interfaces could become fully generative, dynamically adapting in real-time based on user inputs, contexts, and intentions (Deka et al., 2017).
In this work, we study whether an operating system-like interface can be modeled using a neural generative model, rather than a manually programmed system. In doing so, we investigate the modeling and optimization challenges, such as precise cursor control and state tracking. Such a neural interface provides a basis for human-computer interaction research on user-controllable UI generation (e.g., prompt-to-UI personalization (Tolomei et al., 2023), and provides a safe environment in which computer-use agents can be trained and evaluated without issuing real system commands.
To this end, we introduce NeuralOS, a first step toward realizing this vision. NeuralOS simulates an operating system’s graphical interface entirely using deep neural networks. By modeling the OS interface as a generative process, it directly predicts graphical frames from user input events, such as mouse movements, clicks, and keyboard interactions, without manually programmed kernels or applications. Figure 1 illustrates an example sequence generated by NeuralOS, demonstrating realistic cursor movements and window interactions predicted solely from user inputs.
Figure 1: Real image sequence predicted by NeuralOS, illustrating the model’s ability to simulate realistic GUI interactions. The sequence shows key frames as a user (a–c) opens and closes the “Home” folder, followed by (d–f) launches and closes a Doom application that was trained into the model using synthetic demonstrations. Cursor positions are highlighted with red circles. Frames are generated autoregressively, conditioned on previous frames and user inputs.
NeuralOS integrates two complementary architectures, analogous to the separation between OS kernels and desktop rendering programs: a recurrent neural network (Hochreiter & Schmidhuber, 1997) maintains internal computer states (such as open applications, hidden windows, and recent actions), while a diffusion-based convolutional neural renderer generates screen images. We train NeuralOS end-to-end on interaction sequences recorded from Ubuntu XFCE environments, combining randomly generated and realistic AI-generated human-like interactions.
@@
Experiments show that NeuralOS can generate realistic screen sequences, accurately model mouse interactions, and reliably simulate transitions such as application launches. While computational constraints limit its ability to precisely model fine-grained keyboard inputs, NeuralOS represents a step toward neural operating systems that adapt interfaces in real time, potentially enabling users to interact through natural language or gestures rather than fixed menus.
Beyond mimicking an existing operating system, NeuralOS can in principle learn user interfaces from demonstrations even if they are artificially constructed and do not exist in reality. As a proof of concept, we created a Doom application by combining fabricated desktop interactions with VizDoom gameplay recordings. NeuralOS learned to launch, play, and close Doom despite the application never being installed in the underlying system. This illustrates the broader principle that synthetic demonstrations, once distilled into a generative model, become usable user interfaces.
Our code, pretrained models, and an interactive demo are at https://neural-os.com.
@@ 2 Generative Modeling of Operating System Interfaces
Formally, each frame $x_{t}$ is represented as an image tensor $x_{t}\in\mathbb{R}^{H\times W\times C}$, with $H$ and $W$ denoting image height and width, and $C$ the number of color or feature channels. The input event $a_{t}$ at timestep $t$ includes cursor coordinates $(x,y)$, binary indicators for mouse clicks (left or right), and a binary vector indicating which keyboard keys are pressed or released.
The probability distribution of an OS graphical sequence given user inputs can be expressed as:
@@
where $\theta$ represents the parameters of the neural generative model.
Unlike video generation (Singer et al., 2022b; Ho et al., 2022b; OpenAI, 2024), OS simulation must respond instantly to unpredictable user inputs, often causing abrupt changes in the interface, such as when a new application is launched. This contrasts with the smooth, predictable transitions typical in video generation. As a result, the model must maintain accurate and responsive state tracking. Next, we describe the NeuralOS architecture and training strategies designed for these requirements.
[…]
@@ 3 NeuralOS Architecture
NeuralOS employs a hierarchical two-level RNN architecture to track the system state (see Figure 2(b)). Unlike transformers (Vaswani et al., 2017), whose inference complexity increases with context length, the RNN maintains constant complexity per timestep, which is crucial for continuous, long-horizon OS simulation. Existing video-game models (Feng et al., 2024; Oh et al., 2015; Valevski et al., 2024) typically rely on short context windows, which are sufficient because most game states are visually observable from recent frames. In contrast, the RNN state enables NeuralOS to recall previous interactions far in the past (Table 2).
At each timestep $t$, user inputs $a_{t}$ are encoded into embeddings. Specifically, cursor coordinates are discretized screen positions ($a_{t}^{x}$, $a_{t}^{y}$), mouse clicks are binary indicators, and keyboard keys are binary press/release states. Each component is embedded separately and then concatenated:
@@
To ensure that the lower-level LSTM maintains awareness of higher-level contexts, the upper-level LSTM’s output $U_{t}$ is fed back as an input to the lower-level LSTM in the next timestep.
@@
Precise cursor localization is critical for realistic OS interactions. Interactive world and videogame models (Parker-Holder & Fruchter, 2025; Xiang et al., 2025; Alonso et al., 2024; Valevski et al., 2024) typically use a few actions to navigate the world. In contrast, the cursor represents a much larger discrete action space that grows in screen size. This made cursor rendering extremely challenging, where errors in cursor location were often hundreds of pixels as shown in Figure 4(b). To counter this, NeuralOS explicitly encodes cursor positions using a Gaussian spatial map $M_{t}\in\mathbb{R}^{H\times W}$. Instead of using a one-hot cursor position (which can lose precision due to latent resolution constraints), we construct a Gaussian map centered at the cursor’s scaled coordinates:
@@
As demonstrated in Figure 4(b), this spatial encoding is vital for accurate cursor rendering. This map $M_{t}$, combined with LSTM outputs $L_{t}$, $U_{t}$, forms the renderer context $R_{t}\in\mathbb{R}^{H\times W\times C^{\prime}}$:
@@
We concatenate the noisy image with $R_{t}$ as input to the UNet, and then predict the clean image.
@@
Figure 3: Multi-stage training pipeline for NeuralOS. (1) RNN Pretraining: The RNN is pretrained to predict latent frames using a mean squared error (MSE) loss. (2) Joint Training: The pretrained RNN and the diffusion-based renderer are jointly optimized using a standard diffusion loss. (3) Scheduled Sampling: To mitigate error accumulation caused by exposure bias, the most recent input frame is occasionally replaced by a previously generated frame. (4) Context Length Extension: The input context is extended to enable the model to capture long-term dependencies.
@@ 4 Multi-Stage Training Approach
Training NeuralOS is challenging due to ineffective use of RNN outputs, error accumulation during inference, and difficulties in capturing long-term dependencies due to computational constraints. To address these challenges, we take a multi-stage training approach (Figure 3).
@@
Unlike text-to-image diffusion models (Rombach et al., 2022), which use pretrained textual encoders, NeuralOS uses a customized RNN without pretrained checkpoints. Our initial experiments show that direct joint training leads to the renderer ignoring RNN outputs, as indicated by negligible gradient flow into the RNN. The diffusion-based renderer receives two streams of inputs: noisy latent frames and the RNN output; and without proper initialization of the RNN, it relies solely on the noisy image inputs, resulting in an under-trained RNN.
To address this, we first pretrain the RNN. We structure the RNN output $R_{t}\in\mathbb{R}^{H\times W\times C^{\prime}}$ to match the spatial dimensions of the latent frame $x_{t}\in\mathbb{R}^{H\times W\times C}$, but with more channels ($C^{\prime}>C$). The RNN is pretrained to predict the latent frames $x_{t}$ using a mean squared error (MSE) loss:
@@
We jointly optimize the pretrained RNN and the diffusion-based renderer with a standard diffusion loss. The meaningful latent representations learned in RNN pretraining enable the renderer to use the RNN outputs, thus preventing the RNN outputs from being ignored.
@@
During inference, errors accumulate over time and progressively degrade the quality of the generated frames. This issue arises from exposure bias (Ranzato et al., 2015): a model trained exclusively on ground-truth frames becomes overly reliant on perfect inputs and struggles when forced to operate on its own imperfect predictions during inference.
@@
Although NeuralOS can model sequences of arbitrary length, hardware memory limits justify shorter sequences during training, which limits exposure to long-term dependencies. To address this, we introduce a final stage that extends training to longer contexts, following initial training on short context windows for efficiency.
@@ 5 Data Collection
To collect realistic user interactions, we use Anthropic’s Claude-3.5-Sonnet computer-use agent (Anthropic, 2024), which processes screenshots and invokes provided interaction functions. To maximize interaction diversity, we structure the agent’s exploration process around a state-space search tree representing various OS states (see Figure 13 in Appendix J).
@@
In generating these random interactions, we simulate mouse movements, clicks, and keyboard inputs (key presses and releases) stochastically. To improve realism, we introduce several constraints and heuristics iteratively developed through experimentation. Cursor movements are modeled using Bezier curves to emulate natural mouse trajectories (Mortenson, 1999). Double-click events, rare under purely random sampling, are explicitly generated. Additionally, we enforce constraints such as limiting simultaneous key presses and ensuring keys are released only if previously pressed.
Figure 4: (a) Heatmap illustrating predicted vs. ground truth state transitions. Each cell represents the percentage of predictions assigned to a particular predicted cluster (x-axis), given a ground-truth cluster (y-axis). Only the top 16 clusters are displayed here; refer to Figure 11 for the complete heatmap. (b) Comparison of cursor position errors for NeuralOS (with cursor position map), NeuralOS without the cursor position map, and a random baseline. (c) Pixel RMSE of generated frames of using versus not using stage 3 scheduled sampling training.
@@ 6 Experimental Setup
We collected data using 64 parallel Docker containers, each with Ubuntu 20.04 and XFCE desktops at a resolution of $512\times 384$. To simplify the environment, the desktop was limited to four applications: Home, Trash, Terminal, and Firefox. Data consisted of 2K agent-based and 120K random exploration demonstrations, each 30 seconds long at 15 fps, resulting in 12TB of latent data after compression via an autoencoder. The autoencoder reduced the images by a factor of 8 to a latent resolution of $64\times 48$ with 16 channels. Details of the autoencoder are in Appendix I.
Model
The hierarchical RNN has two LSTM modules (each with hidden size 4,096) and a multi-headed attention module (8 heads, 1,024 total dimension). The RNN output is projected to 32 channels and concatenated with the noisy latent frame (16 channels). The UNet uses four resolution levels with channel multipliers of [1, 2, 3, 5], two residual blocks per level, and attention layers at resolutions 8, 4, and 2. It has a base model dimension of 192 and outputs 16 channels. The final model contains 2.2B parameters for the RNN and 263M parameters for the UNet renderer.
@@
NeuralOS was trained using our proposed multi-stage training approach. See Appendix C for full details on hyperparameters. The total data processing and training took approximately 4 months, requiring 17,000 GPU hours on a server with 8 NVIDIA H200 GPUs (141GB memory per GPU) and an additional 6,000 GPU hours on a server with 8 NVIDIA H100 GPUs (80GB memory per GPU). At inference time, we used DDIM sampling (Song et al., 2020a) with 2 steps, achieving an inference speed of 18 fps on a single NVIDIA H100 GPU.